OpenStudy (anonymous):

Write an equation for \[ d \rho \] in terms of \[ \rho \] and \[ dT \] if \[ \rho \] is a function of V and \[ dV = \beta V dT \]

OpenStudy (jamesj):

For what it's worth, are all of these physical quantities of an ideal gas, or perhaps a solid?: rho = density T = temperature V = volume

OpenStudy (anonymous):

Yes that's the idea Can you get \[ d \rho \] from \[d \rho/dt = d \rho/dV*dV/dt\]

OpenStudy (anonymous):

but what is \[ d \rho/dV ?\]

OpenStudy (anonymous):

oops I suppose those little t's above should be T's

OpenStudy (jamesj):

By definition Mass = rho.V and as mass is constant, then d(rho)/dV = d(mass/V)/dV = -mass/V^2

OpenStudy (anonymous):

how would you get that in terms of \[ d \rho, \rho,\] and \[dT \]?

OpenStudy (jamesj):

...so d(rho)/dV = -mass/V^2 = -mass/(mass/rho)^2 = -rho^2/mass Now use the chain rule you have above and the fact that dV/dT = beta.T

OpenStudy (anonymous):

that would work great...but mass isn't specified in this problem.

OpenStudy (jamesj):

no, but you can eliminate again with rho and V if you need to.

OpenStudy (jamesj):

...I was sort of hoping by now you have seen that.

OpenStudy (anonymous):

haha I just did as you were typing.

OpenStudy (anonymous):

wait. so the units check out, but wouldn't that leave \[d \rho/dT = \beta T(- \rho^2/mass)\]

OpenStudy (jamesj):

Right. Now if dV = beta.V dT, then you can solve for V in terms of T. Hence you can substitute into your equation above with V but then write it as V(T). Then the only variables will be T and rho

OpenStudy (anonymous):

you mean integrate?

OpenStudy (jamesj):

i.e. dV/dT = beta.V ==> V = V_0.exp(beta.T)

OpenStudy (anonymous):

yea. ya did. haha

OpenStudy (jamesj):

Ok ... I'm off to bed. Good luck.

OpenStudy (anonymous):

thank you sir